OpenStudy (fantalux):

PLZ HELP! List the two solutions for: 2 |n + 3| = 10. Solve for x: 2(5x + 3) = 7x - 12.

OpenStudy (anonymous):

2|n+3|=10 |n+3|=5 n+3=5 or n+3=-5 n=2 or n=-8

OpenStudy (fantalux):

Thankyou! :)

OpenStudy (anonymous):

2(5x+3)=7x-12 10x+6=7x-12 3x=18 x=6